A pond can be filled with runoff water in 6 hours. It can be filled with well water in 9 hours. How long will it take to fill the pond using both runoff and well water. Thanks :)
Answer by nerdybill(7384)

A pond can be filled with runoff water in 6 hours. It can be filled with well water in 9 hours. How long will it take to fill the pond using both runoff and well water.
.
Let x = time it takes for both to fill pond
then
x(1/6 + 1/9) = 1
multiplying both sides by 54:
x(9 + 6) = 54
15x = 54
x = 54/15
x = 3.6 hours
or
x = 3 hours and 36 minutes